A thin spherical shell of radius R = 20.0 cm has total charge Q = 58.0 nC uniformly distributed on its surface.

Part A

A test particle with charge q = 8.00 nC is initially at a position r = 43.0 cm from the center of the shell. The particle moves to the surface of the shell closest to its initial position. What is the change in potential energy of the test particle as a result of this move?

Part B

What is the potential difference between the final position of the test charge in part A and the initial position of the test charge in part A?

Part C

The test particle now moves from the surface of the charged shell to the center of the shell. What is the change in potential energy of the test particle as a result of this 20.0 cm move?

Part D

What is the potential difference between the center of the shell and the surface of the shell?
Answer #1

(a) Write the equation of Gauss law. E-A= E(4T2) _ E= 4TE 4πε,r Write the formula for the potential E dr V=J - dr 4πε, = 4T

1 20cm for r and 43cmfor, in (2). 47tE Substitute 58.0 nC for q,, 9x10Nm2/C2 for- 1 (58.0nC) (9x109 Nm2/C2 1 AV 20 cm 43 cm

(b) The potential difference between the final position of the test charge in part A and the initial position of the test cha

AV =0 V -_4 4πεJ, Rewrite the above equation. AV. (7) 4πε. 1 and 20cm forr, in (7) Απε, Substitute 58.00 nC for q.. 9x10° N

Hence, the change in potential energy is-1.73x10 J (d) The potential difference betweenthe center of the shell and the surfac